DwightH:

Luke is building a storage building and needs to have a slab of concrete poured. He knows the contractor charges an initial cost of $95 plus an additional $3.50 per square foot of concrete. Which function can be used to determine the cost, in dollars, to pour a concrete slab with an area of x square feet?

DwightH:

A) f(x) = 3.5x + 95 B) f(x) = 9.5x + 3.5 C) f(x) = 95x + 3.5 D) f(x) = 98.5x

dude:

Well usually in these types of questions you want to look at the wording, even if you are not sure what they are talking about When they mention \(\bf per~\#\), it refers to a rate - We know that a rate is always the number next to the variable (x)